# Beyond Pixels: Enhancing LIME with Hierarchical Features and Segmentation Foundation Models

### v4

@@ Abstract
LIME (Local Interpretable Model-agnostic Explanations) is a well-known XAI framework for unraveling decision-making processes in vision machine-learning models. The technique utilizes image segmentation methods to identify fixed regions for calculating feature importance scores as explanations. Therefore, poor segmentation can weaken the explanation and reduce the importance of segments, ultimately affecting the overall clarity of interpretation. To address these challenges, we introduce the DSEG-LIME (Data-Driven Segmentation LIME) framework, featuring: i) a data-driven segmentation for human-recognized feature generation by foundation model integration, and ii) a user steered granularity in the hierarchical segmentation procedure through composition. Our findings demonstrate that DSEG outperforms on several XAI metrics on pre-trained ImageNet models and improves the alignment of explanations with human-recognized concepts.
@@ 1 Introduction
Why should we trust you? The integration of AI-powered services into everyday scenarios, with or without the need for specific domain knowledge, is becoming increasingly common. For instance, consider AI-driven systems that assist in diagnosing diseases based on medical imaging. In such high-stakes scenarios, accuracy and alignment with expert knowledge are paramount. To ensure reliability, stakeholders, frequently seek to evaluate the AI’s performance post-deployment. For example, one might assess whether the AI correctly identifies anomalies in medical scans that could indicate early-stage cancer. The derived question - ”Why should we trust the model?” - directly ties into the utility of Local Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016). LIME seeks to demystify AI decision-making by identifying key features that influence the output of a model, underlying the importance of the Explainable AI (XAI) research domain, particularly when deploying opaque models in real-world scenarios (Barredo Arrieta et al., 2020; Linardatos et al., 2021; Garreau & Mardaoui, 2021).
@@
Figure 1: DSEG vs. other segmentation techniques within LIME. LIME-generated explanations (Ribeiro et al., 2016) for EfficientNetB4 (Tan & Le, 2019) using various segmentation methods: DSEG (ours) with SAM (Kirillov et al., 2023), Quickshift (Hoyer et al., 2019), SLIC (Achanta et al., 2012), Felzenszwalb (Felzenszwalb & Huttenlocher, 2004), and Watershed (Neubert & Protzel, 2014). The top predictions are ‘Acoustic Guitar’ ($p$ = 0.31), ‘Golden Retriever’ ($p$ = 0.24), and ‘Electric Guitar’ ($p$ = 0.07).
Segmentation is key. LIME uses segmentation techniques to identify and generate features to determine the key areas of an image that are critical for classification. However, a challenge emerges when these segmentation methods highlight features that fail to align with identifiable, clear concepts or arbitrarily represent them. This issue is particularly prevalent with conventional segmentation techniques. These methods, often grounded in graph- or clustering-based approaches (Wang et al., 2017), were not initially designed for distinguishing between different objects within images. However, they are the default in LIME’s implementation (Ribeiro et al., 2016).
Ambiguous explanations. The composition of the segmentation has a significant influence on the explanation’s quality (Schallner et al., 2020). Images with a large number of segments frequently experience significant stability issues in LIME, primarily due to the increased number of sampled instances (Section 2). This instability can lead to the generation of two entirely contradictory explanations for the same example, undermining trust not only in LIME’s explanations but also in the reliability of the model being analyzed (Garreau & Mardaoui, 2021; Alvarez-Melis & Jaakkola, 2018; Zhou et al., 2021; Zhao et al., 2020; Tan et al., 2024). Moreover, humans often struggle to interpret the explanations, as the highlighted areas do not align with our intuitive understanding (Molnar et al., 2022; Kim et al., 2022).
This work. In this paper, we address the challenges above by introducing DSEG-LIME (Data-driven Segmentation LIME), an adaptation of the LIME framework for image analysis across domains where specialized knowledge is not required. We replace the conventional segmentation algorithm with foundation models, thereby allowing the end user to employ their model of choice, such as SAM (Kirillov et al., 2023). We often refer to these models as data-driven to emphasise their capability to generate features that more effectively capture human-recognisable concepts, leveraging insights derived from extensive image datasets. Given the great segmentation ability of such models, we implement a compositional object structure, adapting LIME’s feature generation with a novel hierarchical segmentation. This adaptation provides flexibility in the granularity of concepts, allowing users to specify the detail of LIME’s explanation, viewing a car as a whole or in parts like doors and windshields. This approach breaks down broad categorizations, enabling independent evaluation of each sub-concept. Figure 1 demonstrates the motivation mentioned above by employing LIME, which generates explanations using various segmentation techniques, specifically focusing on an image of a dog playing the guitar. In this context, DSEG excels by more clearly highlighting features that align with human-recognizable concepts, distinguishing it from other methods.
Contribution. The key contributions of our paper are summarized as follows: (i) We present DSEG-LIME, an enhanced version of the LIME framework for image analysis, leveraging foundation models to improve image segmentation. (ii) DSEG extends LIME by incorporating compositional object structures, enabling hierarchical segmentation that offers users adjustable feature granularity. (iii) We rigorously evaluate our approach with other segmentation methods and LIME enhancements across multiple pre-trained image classification models. Our evaluation includes a user study for qualitative insights and distinguishes between explaining (quantitative) and interpreting (qualitative) aspects. We acknowledge that explanations considered intuitive by users may not always reflect the AI model’s operational logic, which can diverge from human perception (Molnar et al., 2022; Freiesleben & König, 2023). To address this, we complement our evaluation with several quantitative performance metrics widely used in XAI research (Nauta et al., 2023).
@@ 2 Related work
Region-based perturbation XAI techniques. LIME is among several techniques designed to explain black box models through image perturbation. Fong & Vedaldi (2017) introduced a meta-predictor framework that identifies critical regions via saliency maps. Subsequently, Fong et al. (2019) developed the concept of extremal perturbations to address previous methods’ limitations. Additionally, Kapishnikov et al. (2019) advanced an integrated-gradient, region-based attribution approach for more precise model explanations. More recently, Escudero-Viñolo et al. (2023) have highlighted the constraints of perturbation-based explanations, advocating for the integration of semantic segmentation to enhance image interpretation.
@@
Instability of LIME. The XAI community widely recognizes the instability in LIME’s explanations, which stems from LIME’s design (Alvarez-Melis & Jaakkola, 2018; Zhou et al., 2021; Zhao et al., 2020; Tan et al., 2024). Alvarez-Melis & Jaakkola (2018) handled this issue by showing the instability of various XAI techniques when slightly modifying the instance to be explained. A direct improvement is Stabilized-LIME (SLIME) proposed by Zhou et al. (2021) based on the central limit theorem to approximate the number of perturbations needed in the data sampling approach to guarantee improved explanation stability. Zhao et al. (2020) improved stability by exploiting prior knowledge and using Bayesian reasoning - BayLIME. GLIME (Tan et al., 2024) addressed this issue by employing an improved local and unbiased data sampling strategy, resulting in explanations with higher fidelity - similar to the work by Rashid et al. (2024). Recent advancements include Stabilized LIME for Consistent Explanations (SLICE) (Bora et al., 2024), which improves LIME through a novel feature selection mechanism that removes spurious superpixels and introduces an adaptive perturbation approach for generating neighbourhood samples. Another hierarchical-based variation, DLIME (Zafar & Khan, 2021), utilizes agglomerative hierarchical clustering to organize training data, focusing primarily on tabular datasets. In contrast, DSEG-LIME extends this concept to images by leveraging the hierarchical structure of image segments.
@@
Segmentation influence on explanation. The segmentation algorithm utilized to sample data around the instance $\mathbf{x}$ strongly influences its explanation. It directly affects the stability of LIME itself, as suggested by Ng et al. (2022). This behaviour is in line with the investigation by Schallner et al. (2020) that examined the influence of different segmentation techniques in the medical domain, showing that the quality of the explanation depends on the underlying feature generation process. Blücher et al. (2024) explored how occlusion and sampling strategies affect model explanations when integrated with segmentation techniques for XAI, including LRP (Layer-Wise Relevance Propagation) (Montavon et al., 2019) and SHAP (Lundberg & Lee, 2017). Their study highlights how different strategies provide unique explanations while evaluating the SAM technique in image segmentation. Sun et al. (2023) used SAM within the SHAP framework to provide conceptually driven explanations, which we discuss in Section B.4 and further supports our approach.
@@
Segmentation hierarchy. Li et al. (2022) aimed to simulate the way humans structure segments hierarchically and introduced a framework called Hierarchical Semantic Segmentation Networks (HSSN), which approaches segmentation through a pixel-wise multi-label classification task. HIPPIE (HIerarchical oPen-vocabulary, and unIvErsal segmentation), proposed by Wang et al. (2023), extended hierarchical segmentation by merging text and image data multimodally. It processes inputs through decoders to extract and then fuse visual and text features into enhanced representations.
@@ 3 Foundations of LIME
LIME is a prominent XAI framework to explain a neural network $f$ in a model-agnostic and instance-specific (local) manner. It applies to various modalities, including images, text, and tabular data (Ribeiro et al., 2016). In the following, we will briefly review LIME’s algorithm for treating images, also visualized in Figure 2 next to our adaption with DSEG.
Figure 2: DSEG Pipeline in LIME. LIME image analysis pipeline with DSEG integrated into feature generation. Dashed lines indicate the option to apply DSEG.
Notation. We consider a scenario where we deal with imagery data. Let $\mathbf{x}\in\mathcal{X}$ represent an image within a set of images, and let $\mathbf{y}\in\mathcal{Y}$ denote its corresponding label in the output space with logits $\mathcal{Y}\subseteq\mathbb{R}$ indicating the labels in $\mathcal{Y}$. We denote the neural network we want to explain by $f:\mathcal{X}\rightarrow\mathcal{Y}$. The network accepts an input $\mathbf{x}$ and produces an output in $\mathcal{Y}$, which signifies the probability $p$ of the instance being classified into a specific class.
Feature generation. The technique involves training a local, interpretable surrogate model $g\in G$, where $G$ is a class of interpretable models, such as linear models or decision trees, which approximates $f$’s behavior around an instance $\mathbf{x}$ (Ribeiro et al., 2016). This instance needs to be transformed into a set of features that can be used by $g$ to compute the importance score of its features. In the domain of imagery data, segmentation algorithms segment $\mathbf{x}$ into a set of superpixels $s_{0}...s_{d}\in\mathcal{S}^{D}$, done by conventional techniques (Hoyer et al., 2019). We use these superpixels as the features for which we calculate their importance score. This step illustrates the motivation of Figure 1, which underpins the quality of features affecting the explanatory quality of LIME.
Sample generation. For sample generation, the algorithm manipulates superpixels by toggling them randomly. Specifically, each superpixel $s_{i}$ is assigned a binary state, indicating this feature’s visibility in a perturbed sample $\mathbf{z}$. The presence (1) or absence (0) of these features is represented in a binary vector $\mathbf{z^{\prime}_{i}}$, where the $i$-th element corresponds to the state of the $i$-th superpixel in $\mathbf{z}$. When a feature $s_{i}$ is absent (i.e., $s_{i}=0$), its pixel values in $\mathbf{z}$ are altered. This alteration typically involves replacing the original pixel values with a non-information holding value, such as the mean pixel value of the image or a predefined value (e.g., black pixels) (Ribeiro et al., 2016; Tan et al., 2024).
@@
In this equation, $\mathbf{z}$, and $\mathbf{z^{\prime}}$ are sampled instances from the perturbed dataset $\mathcal{Z}$, and $g$ is the interpretable model being learned (Ribeiro et al., 2016; Tan et al., 2024). The interpretability of the model is derived primarily from the coefficients of $g$. These coefficients quantify the influence of each feature on the model’s prediction, with each coefficient’s magnitude and direction (positive or negative) indicating the feature’s relative importance and effect.
DSEG breaks up this fixed structure and allows a repeated calculation of features based on the concept hierarchy and user preferences.
@@ 4 DSEG-LIME
In this section, we will present DSEG-LIME’s two contributions: first, the substitution of traditional feature generation with a data-driven segmentation approach (Section 4.1), and second, the establishment of a hierarchical structure that organizes segments in a compositional manner (Section 4.2).
@@ 4.1 Data-driven segmentation integration
DSEG-LIME improves the LIME feature generation phase by incorporating data-driven segmentation models, outperforming conventional graph- or cluster-based segmentation techniques in creating recognizable image segments across various domains. Specifically, our approach mainly uses SAM (Segment Anything) (Kirillov et al., 2023) due to its remarkable capability to segment images in diverse areas. However, as the appendix shows, it can also be applied to other segmentation models. Figure 2 illustrates the integration of DSEG into the LIME framework, as outlined in Section 3. Specifically, DSEG impacts the feature generation phase, influencing the creation of superpixels/features $\mathcal{S}^{D}$, and subsequently affecting the binary vector $\mathbf{z^{\prime}}$ and the feature vector $\mathbf{z}$. This modification directly impacts the loss function in Equation 1 and the proximity metric for perturbed instances, leading to an improved approximation of the interpretable model $g$, which is used to explain the behaviour of the original model $f$ for a given instance $\mathbf{x}$. However, the effect of DSEG aligns with that of other segmentation methods like SLIC, as the surrogate model primarily leverages the resulting segments without incorporating additional elements from the segmentation foundation models underlying DSEG. This argument is further substantiated in the discussion of our experimental results.

### v5

@@ Abstract
LIME (Local Interpretable Model-agnostic Explanations) is a well-known XAI framework for unraveling decision-making processes in vision machine-learning models. The technique utilizes image segmentation methods to identify fixed regions for calculating feature importance scores as explanations. Therefore, poor segmentation can weaken the explanation and reduce the importance of segments, ultimately affecting the overall clarity of interpretation. To address these challenges, we introduce the DSEG-LIME (Data-Driven Segmentation LIME) framework, featuring: i) a data-driven segmentation for human-recognized feature generation by foundation model integration, and ii) a user-steered granularity in the hierarchical segmentation procedure through composition. Our findings demonstrate that DSEG outperforms on several XAI metrics on pre-trained ImageNet models and improves the alignment of explanations with human-recognized concepts.
[…]
3235
@@ 1 Introduction
Why should we trust you? The integration of AI-powered services into everyday scenarios, with or without the need for specific domain knowledge, is becoming increasingly common. Consider autonomous driving or surveillance, where accurate visual recognition of images or video streams is critical. In such high-stakes applications, precision and alignment with expert judgment and safety standards are paramount. To ensure reliability and safety, stakeholders frequently evaluate AI performance after deployment. For example, one might assess whether an AI correctly identifies pedestrians or hazardous objects in real-time footage, accurately detecting potential threats to avoid accidents or breaches in security. The derived question - "Why should we trust the model?" - is a direct reference to Local Interpretable Model-agnostic Explanations (LIME) [49]. LIME seeks to demystify AI decision-making by identifying key features that influence the output of a model, underlying the importance of the Explainable AI (XAI) research domain, particularly when deploying opaque models in real-world scenarios [5, 20, 32].
@@
Figure 1: DSEG vs. other segmentation techniques within LIME. LIME-generated explanations [49] for EfficientNetB4 [53] using various segmentation methods: DSEG (ours) with SAM [28], Quickshift [23], SLIC [1], Felzenszwalb [16], and Watershed [41]. The top predictions are ‘Acoustic Guitar’ ($p$ = 0.31), ‘Golden Retriever’ ($p$ = 0.24), and ‘Electric Guitar’ ($p$ = 0.07).
Ambiguous explanations. LIME relies on conventional superpixel segmentation (e.g., graph‑ or clustering‑based methods [55]) to isolate image regions for classification explanations. However, these regions often lack semantic coherence or compositional structure, such as distinguishing car doors, wheels, or headlights, which undermines both clarity and relevance. Because LIME defaults to these arbitrary segmentation methods [49], highlighted areas often conflict with human intuition [27, 38]. Moreover, segmentation granularity has a dramatic impact on explanation quality [50]: excessive segments introduce instability, causing LIME to produce mutually contradictory explanations for the same image and eroding trust in both the explanation mechanism and the model itself [4, 20, 54, 60, 61].
This work. We introduce DSEG-LIME (Data-driven Segmentation LIME), an enhanced LIME framework that replaces traditional superpixel algorithms with flexible foundation-model segmenters. Users can plug in general-purpose models such as SAM (Segment Anything) [28] or domain-adapted variants through SAM adapters [11], producing semantically coherent regions that mirror human concepts. Building on these powerful segmenters, we propose a hierarchical segmentation scheme that encodes a compositional object structure. This lets users dial in the granularity of explanations, viewing, for example, a car as a single unit or decomposed into doors, wheels, or headlights. By enabling concept-level analysis at multiple scales, DSEG-LIME delivers clearer, more relevant feature highlights and allows independent evaluation of each subconcept. As shown in Figure˜1, when applied to an image of a dog playing guitar, DSEG-LIME distinctly pinpoints meaningful parts (e.g., the dog’s snout, the guitar’s body), demonstrating superior alignment with human intuition compared to existing LIME variants.
Key contributions. We advance explainable image analysis by:
  • Introduce DSEG‑LIME, the first adaptation of LIME to harness foundation segmentation models, enabling semantically coherent region proposals.
  • Develop a hierarchical segmentation framework embedding a compositional object structure, empowering users with fine‑grained control over explanation granularity—from whole objects to individual parts (e.g., doors, wheels, headlights).
  • Evaluate DSEG‑LIME against conventional segmentation and LIME variants on multiple pretrained classifiers, employing quantitative fidelity metrics [40] and a qualitative user study to assess both explanatory accuracy and human interpretability, while highlighting gaps between human intuition and model reasoning [38, 19].
@@ 2 Related work
Region-based perturbation XAI techniques. LIME is among several techniques designed to explain black box models through image perturbation. Fong and Vedaldi [18] introduced a meta-predictor framework that identifies critical regions via saliency maps. Subsequently, Fong et al. [17] developed the concept of extremal perturbations to address previous methods’ limitations. Additionally, Kapishnikov et al. [24] advanced an integrated-gradient, region-based attribution approach for more precise model explanations. More recently, Escudero-Viñolo et al. [15] highlight the constraints of perturbation-based explanations, advocating for the integration of semantic segmentation to enhance image interpretation.
@@
Instability of LIME. The XAI community widely recognizes the instability in LIME’s explanations, which stems from LIME’s design [4, 54, 60, 61]. Alvarez-Melis and Jaakkola [4] handle this issue by showing the instability of various XAI techniques when slightly modifying the instance to be explained. A direct improvement is Stabilized-LIME (SLIME) proposed by Zhou et al. [61] based on the central limit theorem to approximate the number of perturbations needed in the data sampling approach to guarantee improved explanation stability. Zhao et al. [60] improve stability by exploiting prior knowledge and using Bayesian reasoning - BayLIME. GLIME [54] addresses this issue by employing an improved local and unbiased data sampling strategy, resulting in explanations with higher fidelity - similar to the work by Rashid et al. [47]. Recent advancements include Stabilized LIME for Consistent Explanations (SLICE) [7], which improves LIME through a novel feature selection mechanism that removes spurious superpixels and introduces an adaptive perturbation approach for generating neighborhood samples. Another hierarchical-based variation, DLIME [57], utilizes agglomerative hierarchical clustering to organize training data, focusing on tabular datasets. In contrast, DSEG-LIME extends this concept to images by leveraging the hierarchical structure of image segments.
@@
Segmentation influence on explanation. The segmentation algorithm utilized to sample data around the instance $\mathbf{x}$ strongly influences its explanation. It directly affects the stability of LIME itself, as suggested by Ng et al. [42]. This behavior is in line with the investigation by Schallner et al. [50] that examines the influence of different segmentation techniques in the medical domain, showing that the quality of the explanation depends on the underlying feature generation process. Bluecher et al. [6] explore how occlusion and sampling strategies affect model explanations when integrated with segmentation techniques for XAI, including LRP (Layer-Wise Relevance Propagation) [39] and SHAP [34]. Their study highlights how different strategies provide unique explanations while evaluating the SAM technique in image segmentation. Sun et al. [52] use SAM within the SHAP framework to provide conceptually driven explanations, which we discuss in Appendix B.4 [29]. This work reinforces the strategy of employing deeper models to elucidate simpler ones by seamlessly integrating these approaches into widely adopted XAI frameworks.
@@
Segmentation hierarchy. Hierarchical Segment Grouping (HSG) [25] performs unsupervised hierarchical semantic segmentation by leveraging multiview cosegmentation and clustering transformers to enforce spatial and semantic consistency across images. The work of Li et al. [30] aims to simulate the way humans structure segments hierarchically and introduce a framework called Hierarchical Semantic Segmentation Networks (HSSN), which approaches segmentation through a pixel-wise multi-label classification task. HIPPIE (HIerarchical oPen-vocabulary, and unIvErsal segmentation), proposed by Wang et al. [56], extends hierarchical segmentation by merging text and image data multimodally. It processes inputs through decoders to extract and then fuse visual and text features into enhanced representations. While numerous hierarchical segmentation methods exist, their specialized, end‑to‑end pipelines cannot accommodate external foundation models—making them incompatible with model‑agnostic explainers like LIME. To bridge this gap, we introduce a lightweight, compositional hierarchy within LIME that supports plug‑and‑play integration of any foundation or domain‑specific segmenter.
@@ 3 Foundations of LIME
LIME is a prominent XAI framework to explain a neural network $f$ in a model-agnostic and instance-specific (local) manner. It applies to various modalities, such as images or text [49]. In the following, we will briefly review LIME’s algorithm for treating images, also visualized in Figure˜2 next to our adaptation with DSEG.
Figure 2: DSEG Pipeline in LIME. LIME image analysis pipeline with DSEG integrated into feature generation. Dashed lines indicate the option to apply DSEG. The frozen segmentation foundation model can be freely chosen by the user.
Notation. We consider the scenario in which we deal with image data. Let $\mathbf{x}\in\mathcal{X}$ represent an image within a set of images, and let $\mathbf{y}\in\mathcal{Y}$ denote its corresponding label in the output space with logits $\mathcal{Y}\subseteq\mathbb{R}$ indicating the labels in $\mathcal{Y}$. We denote the neural network we want to explain by $f:\mathcal{X}\rightarrow\mathcal{Y}$. This network functions by accepting an input $\mathbf{x}$ and producing an output in $\mathcal{Y}$, which signifies the probability $p$ of the instance being classified into a specific class.
Feature generation. The technique involves training a local, interpretable surrogate model $g\in G$, where $G$ is a class of interpretable models, such as linear models or decision trees, which approximates $f$’s behavior around an instance $\mathbf{x}$ [49]. This instance needs to be transformed into a set of features that can be used by $g$ to compute the importance score of its features. In the domain of imagery data, segmentation algorithms segment $\mathbf{x}$ into a set of superpixels $s_{0}...s_{d}\in\mathcal{S}^{D}$, done by conventional techniques [23]. We use these superpixels as the features for which we calculate their importance score. This step illustrates the motivation of Figure˜1, which underpins the quality of features affecting the explanatory quality of LIME.
Sample generation. For sample generation, the algorithm manipulates superpixels by toggling them randomly. Specifically, each superpixel $s_{i}$ is assigned a binary state, indicating this feature’s visibility in a perturbed sample $\mathbf{z}$. The presence (1) or absence (0) of these features is represented in a binary vector $\mathbf{z^{\prime}_{i}}$, where the $i$-th element corresponds to the state of the $i$-th superpixel in $\mathbf{z}$. When a feature $s_{i}$ is absent (i.e., $s_{i}=0$), its pixel values in $\mathbf{z}$ are altered. This alteration typically involves replacing the original pixel values with a non-information holding value, such as the mean pixel value of the image or a predefined value (e.g., black pixels) [49, 54].
@@
In this equation, $\mathbf{z}$, and $\mathbf{z^{\prime}}$ are sampled instances of the perturbed data set $\mathcal{Z}$, and $g$ is the interpretable model being learned [49, 54]. Interpretability is derived primarily from the coefficients of $g$. These coefficients quantify the influence of each feature on the model’s prediction, with each coefficient’s magnitude and direction (positive or negative) indicating the feature’s relative importance. DSEG breaks up this fixed structure and allows for a repeated calculation of features based on the concept hierarchy and user preferences.
@@ 4 DSEG-LIME
In this section, we will present DSEG-LIME’s two contributions: first, the substitution of traditional feature generation with a data-driven segmentation approach (Section˜4.1), and second, the establishment of a hierarchical structure that organizes segments in a compositional manner (Section˜4.2).
@@ 4.1 Data-driven segmentation integration
DSEG-LIME expands the LIME feature generation phase by incorporating data-driven segmentation models, outperforming conventional graph- or cluster-based segmentation techniques in creating recognizable image segments across various domains. For the remainder of the paper, we mainly use SAM [28] due to its remarkable capability to segment images in diverse areas. However, in the appendix, we show that it can also be applied to other segmentation models. Figure˜2 illustrates DSEG’s integration into the LIME framework of Section˜3. The input image is first passed to a segmentation foundation model, which outputs a set of segments. Importantly, we only use the generated segments for the explanation process and do not incorporate any additional information from the foundation model. By redefining the feature basis, DSEG alters the loss in Equation˜1 and the proximity metric for sampling, often producing a surrogate $g$ that more faithfully approximates the original model $f$ at instance $\mathbf{x}$. However, the effect of DSEG is comparable to other segmentation methods such as SLIC, since the surrogate model relies solely on the resulting segments and does not incorporate any additional information from the underlying segmentation foundation models.